Hello I need help on this one problem.
How do you write a quadratic function in standard form for the parabola that goes through the points of (-6,-2), (-4,-2), (-3,4).

Answer by scott8148(6628)

use the general quadratic and substitute the points to find the coefficients

ax^2 + bx + c = 0

(-6,-2) ___ a(-6)^2 + b(-6) + c = -2 ___ 36a - 6b + c = -2

(-4,-2) ___ a(-4)^2 + b(-4) + c = -2 ___ 16a - 4b + c = -2

(-3,4) ___ a(-3)^2 + b(-3) + c = 4 ___ 9a - 3b + c = 4

solve the system of equations to find a, b, and c